# Visual Representation Learning Guided by Multi-modal Prior Knowledge

### v1

@@ Abstract
Despite the remarkable success of deep neural networks (DNNs) in computer vision, they fail to remain high-performing when facing distribution shifts between training and testing data. In this paper, we propose Knowledge-Guided Visual representation learning (KGV) – a distribution-based learning approach leveraging multi-modal prior knowledge – to improve generalization under distribution shift. We use prior knowledge from two distinct modalities: 1) a knowledge graph (KG) with hierarchical and association relationships; and 2) generated synthetic images of visual elements semantically represented in the KG. The respective embeddings are generated from the given modalities in a common latent space, i.e., visual embeddings from original and synthetic images as well as knowledge graph embeddings (KGEs). These embeddings are aligned via a novel variant of translation-based KGE methods, where the node and relation embeddings of the KG are modeled as Gaussian distributions and translations respectively. We claim that incorporating multi-model prior knowledge enables more regularized learning of image representations. Thus, the models are able to better generalize across different data distributions. We evaluate KGV on different image classification tasks with major or minor distribution shifts, namely road sign classification across datasets from Germany, China, and Russia, image classification with the mini-ImageNet dataset and its variants, as well as the DVM-CAR dataset. The results demonstrate that KGV consistently exhibits higher accuracy and data efficiency than the baselines across all experiments.
@@ Introduction
Deep neural networks (DNNs) have demonstrated outstanding performance in pattern recognition for computer vision (CV) tasks. They achieved remarkable success by leveraging vast amounts of data and computational power to solve the image classification task. However, these models often struggle with generalization, especially when the test data distribution deviates from the training data distribution (Liang, He, and Tan 2023). This happens as DNNs typically assume the training and test data distribution are identical, and they tend to overfit to the training domain. The situation worsens under low data regimes (Adadi 2021), as DNNs are more prone to overfitting when insufficient training data exists. For example, a road sign perception system trained with road signs from one country will deteriorate classification accuracy when applied in a different country. This is due to the fact that visual elements of the signs vary between countries, as illustrated in Figure 1a). For instance, in Germany, warning signs have a white background, whereas in China, they use a yellow background. Similarly, stop signs display varying text or fonts across different countries.
Refer to caption(a)Germany:China:Russia:Refer to caption(b)TriangleRedWhiteExclamationhasShapehasBorderColorhasBackgroundColorhasSignLegendRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRoad Signs(c)MandatoryWarning Danger
Animal
Danger
Sign
DE CN Triangle Exclamation
Figure 1: (a) Examples of distribution shift of road signs in different countries. (b) Decomposition of visual elements w.r.t. the road sign. (c) Abstract representation of hierarchical and association relations.
@@
First, we capture information about the structures of objects in a KG (Hogan et al. 2021) to inform and enhance the learning process. For instance, warning signs are composed of a white triangle with a red border plus one pictographic element (cf. Figure 1b). Road signs in different countries share the same elements, such as the shape or the pictographic element, or both. An abstract representation of such a KG is demonstrated in Figure 1c. For this, we developed a translation-based KGE method to map the nodes of the KG to embeddings in the latent space. An object category (e.g., ‘danger sign’) is naturally represented by its set of image vectors in the latent space. To align image and object categories, we represent the latter by a Gaussian distribution in the common latent space. Thus, the embeddings of object category elements and their relations to object categories indirectly influence the formation of categories in the latent space.
Second, we generate and use synthetic images of object category elements associated with KG as visual prior knowledge to augment the dataset. For instance, in the road sign domain, these elements are shape, color, and legend within the sign. These images are considered as visual prior knowledge of the object categories defined in the KG. Therefore, in this paper, the prior knowledge we used has two distinct modalities: 1) the factual knowledge across different data distributions represented in a KG, and 2) generated synthetic images of object category elements semantically represented in the KG.
We evaluate our model’s performance under data distribution shifts and low data regimes in the context of image classification. To represent the prior knowledge, we construct three targeted KGs for the domains of Road Sign Recognition, Car Recognition, and ImageNet. Our experimental results show that our KGV method outperforms the image classification baselines in all experiments. Additionally, we found that our method is more data-efficient. Our main contributions are summarized as follows:
@@
  • We evaluated our KGV method on various image classification tasks and datasets, namely road sign datasets, mini-ImageNet and its variants, as well as the DVM-CAR datasets. The results illustrate a notable improvement under distribution shift: 4.4% and 4.1% on average in the ImageNet and Road Sign Recognition domains. Moreover, our KGV method improves the SOTA results on the DVM-CAR dataset and shows significant improvements on data efficiency in low data regimes.
@@
Including information from KGs can enhance model’s ability in visual or textual tasks (Monka, Halilaj, and Rettinger 2022b). Some approaches (Annervaz, Chowdhury, and Dukkipati 2018; Bosselut et al. 2019; Liu et al. 2020) leverage KGs to enhance language models with better reasoning ability. In the field of computer vision, KGs also play an important role. Li et al. (2023a) introduce a KG that includes inter-class relationships to the visual-language model, yielding a more effective classifier for downstream tasks. Some approaches  (Zareian, Karaman, and Chang 2020; Yu et al. 2021) build scene graphs to enhance their models’ performance. There also exist some approaches use extra information in the KG to perform zero-shot learning tasks (Wang, Ye, and Gupta 2018; Kampffmeyer et al. 2019; Roy et al. 2020) or few-shot learning tasks (Li et al. 2023c; Tseng et al. 2020).
@@
Our approach is designed end-to-end by adding the regularization loss and cross-entropy loss together to train the network. The regularization loss aligns the KG embedding and the visual embedding. The cross-entropy loss is used for the image classification. The architecture of KGV is depicted in Figure 2. Our model consists of three phases: the knowledge modeling phase, the training phase, and the inference phase. Here, we take the road sign recognition domain as an example to describe our model.
@@
1. Knowledge Modelling PhaseColorsRefer to captionRefer to captionRefer to captionShapesRefer to captionRefer to captionRefer to captionSign LegendsRefer to captionRefer to captionRefer to captionRoad SignsRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption+Road SignsMandatoryWarning Danger
@@
Animal
Danger
Sign
DECNSign LegendsExclamationAnimalShapesTriangleCirclehasSignLegendOfhasShapeOfRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionimage instancessubClassOfNode Embeddings Relation Embeddings 3. Inference PhaseRefer to caption Image Encoder $z^{I}$ Decoder
@@
Each image is provided as input into the image encoder to obtain the respective image embedding $\boldsymbol{z^{I}}\in\mathbb{R}^{d}$ in the latent space, where $d$ is the size of the dimension. The relations from KG connecting object categories, their elements and their respective images are encoded and stored in the relation lookup table, which is represented as a list of embeddings $[\boldsymbol{z_{0}^{r}},\boldsymbol{z_{1}^{r}},...,\boldsymbol{z_{N_{r}}^{r}}]$, where each $\boldsymbol{z_{i}^{r}}$ is the $i$-th relation vector embedding with the size of $d$. Note that the 0-th relation represents the inclusion relation demonstrated with purple dashed arrows as illustrated in Figure 2. The hierarchy of road signs is preserved in the KG. For instance, the images of the danger signs in Germany belong to the German danger sign category, and they also belong to the danger sign category, warning sign category, and road sign category. We use $\boldsymbol{z_{0}^{r}}$ to represent all these inclusion relations. In terms of the other $N_{r}$ relations, if one node $h$ has one relation with another node $t$, then all the sub-category nodes of $h$ have such relation with the node $t$. For example, all warning signs have a triangle shape, including their sub-categories like danger signs and danger animal signs. The node lookup table is represented as a list of embeddings $[\boldsymbol{z_{1}^{o}},\boldsymbol{z_{2}^{o}},...,\boldsymbol{z_{N_{o}}^{o}}]$, where each $\boldsymbol{z_{j}^{o}}$ denotes the $j$-th embedding. $N_{o}$ is the total number of nodes in the KG. The node embeddings are considered as Gaussian embeddings $\boldsymbol{z_{j}^{o}}=(\boldsymbol{\mu_{j}},\boldsymbol{\Sigma_{j}})$, where $\boldsymbol{\mu_{j}}$ and $\boldsymbol{\Sigma_{j}}$ are the mean and variance of the Gaussian distribution. Note that each node, their respective relations, and images are embedded in the same latent space.
@@
where $\mathcal{N}(\boldsymbol{z^{I}}+\boldsymbol{z_{i}^{r}};\boldsymbol{\mu_{j}},\boldsymbol{\Sigma_{j}})$ denotes the possibility density of the vector $\boldsymbol{z^{I}}+\boldsymbol{z_{i}^{r}}$ under the Gaussian distribution with parameters of $\boldsymbol{\mu_{j}}$ and $\boldsymbol{\Sigma_{j}}$.
@@
where M is the mask defined in Equation (1). $\textbf{S}\in R^{(N_{r}+1)\times N_{o}}$ and $\textbf{S}_{ij}=-\log(s(\boldsymbol{z^{I}},\boldsymbol{z_{i}^{r}},\boldsymbol{z_{j}^{o}}))$, following the negative likelihood form. $\epsilon$ is the threshold score difference between positive and negative triplets. Finally, the total loss is defined by:
@@
In this section, we first explain the datasets and the respective knowledge graphs used in our experiments. Then, we analyze the results of supervised image classification experiments under varying conditions, including data distribution shifts, low-data regimes, and few-shot learning scenarios. The experimental details are listed in the appendix.
@@
The constructed knowledge graph contains 202 nodes and 5 relations. The 5 relations we include are ‘has the background color of’, ‘has the border color of’, ‘has the sign legend of’, ‘has the shape of’, ‘instance/subclass of’.
@@
We leverage WordNet (Miller 1995; Fellbaum 2010), which is a lexical database containing nouns, verbs, adjectives, and adverbs of the English language structured into respective synsets and create the subset of Wordnet by extracting the respective synsets of each label from the mini-ImageNet dataset. These synsets are grouped based on the lexical domain they pertain to, e.g., animal, artifact, or food; thus, we get hierarchical relations between the classes in the mini-ImageNet from the WordNet. We also include two association relations: ‘has color’ and ‘has sketch’. We use the generated pure color images in the domain of road signs and the sketch images from ImageNet-Sketch. Overall, it contains 234 nodes and 3 relations: ‘subclass/instance of’, ‘has color’ and ‘has sketch’.
@@ Results
We conducted a number of supervised image classification experiments under major and minor distribution shifts and under low data regimes. Additionally, to evaluate the model’s adaptability to new data distributions, we also conducted few-shot learning experiments using data with altered distributions.
@@
Model GTSRB (43) CTSD (25) RTSD (36)
ResNet50 98.9% (+0.4%) 71.5% (+3.3%) 72.8% (+6.5%)
DGP 99.0 %(+0.3%) 72.0% (+2.8%) 73.5% (+5.8%)
GCNZ 99.1% (+0.2%) 71.1% (+3.7%) 74.9% (+4.4%)
KGV (ours) 99.3% 74.8% 79.3%
@@
KGV 99.1% 73.5% 78.1%
Table 1: Road Sign Classification Accuracy. Evaluation results with the test data of GTSRB and the full data of CTSD and RTSD. ResNet50+ refers to the ResNet50 model trained with additional synthetic images. KGV- denotes KGV trained without synthetic images. KGV uses vector instead of Gaussian embeddings to represent nodes of KG.
@@
We train our KGV model with the GTSRB dataset plus generated synthetic images. The trained models are then evaluated with test data of GTSRB with 43 classes, the full data of CTSD with 25 shared classes, and the full data of RTSD with 36 shared classes. The shared classes here indicate those classes that both datasets contain. As shown in Table 1, the accuracy drops when distribution shift cases, i.e., from Germany to China and Russia. By incorporating the multi-modal prior knowledge, our KGV reaches an obvious improvement compared to other baselines, such as DGP (Kampffmeyer et al. 2019) and GCNZ (Wang, Ye, and Gupta 2018). These two are ZSL baselines, which leverage additional information from KGs.
Further, we conducted other ablation studies to evaluate the impact of the type of prior knowledge in the final results. ‘ResNet50+’ refers to the model trained with additional synthetic images and ‘KGV-’ denotes the model trained without synthetic images. ‘KGV’ is the ablated version of KGV that employs vector embeddings instead of Gaussian embeddings for representing nodes in the KG. As the experiments indicate, our KGV model surpasses the ResNet50 baseline by 3.3% in CTSD and 6.5% in RTSD. We also observe that the inclusion of synthetic images featuring elements of object categories, such as colors and shapes, enhances the generalization ability of the KGV models across various data distributions. On the other hand, it does not have a significant effect on the ResNet50’s performance. The precision and recall for a subset of classes in CTSD are presented in Table 2. As the table illustrates, the ResNet50 fails to classify specific classes, including danger signs, pedestrian crossing, etc. In contrast, the KGV models exhibit significant improvements in both precision and recall.
Datasets Samples ResNet50 KGNN ResNet50+ KGV
Mini-ImageNet [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 62.6% +4.2% 66.2% +0.6% 59.6% +7.2% 66.8%
ImageNetV2 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 50.5% +2.2% 52.5% +0.2% 48.0% +4.7% 52.7%
ImageNet-R [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 25.5% +15.9% 30.0% +11.4% 36.9% +4.5% 41.4%
ImageNet-A [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 4.7% +2.1% 5.3% +1.5% 5.0% +1.8% 6.8%
@@
We also evaluate our KGV method in the ImageNet domain. The baselines we choose are ResNet50 and KGNN (Monka, Halilaj, and Rettinger 2022a), which represents another approach to integrate prior knowledge into learning pipeline through pre-training. The methods ResNet50+ and KGV are trained with mini-ImageNet, ImageNet-Sketch, and pure color images while the methods KGNN and ResNet50 are trained without any visual priors like sketches and color images. Our model outperforms the baseline in all these datasets. As Table 3 shows, these baselines fail, particularly when the distribution shift increases (from Mini-ImageNet to ImageNetV2, ImageNet-R, and ImageNet-A). The KGV model suffers less when data distribution changes by including the multi-modal prior knowledge. We also observe that incorporating visual priors without the guidance of symbolic knowledge from the KG may negatively impact performance in scenarios where the difference between training and test distributions is minimal. ResNet50’s performance decreases from 62.6% to 59.6% on mini-ImageNet and from 50.5% to 48.0% on ImageNetV2.
Model SimCLR MMCL SupCon ResNet50 KGV (Ours)
ACC 88.8% +7.5% 94.4% +1.9% 93.8% +2.5% 94.3% +2.0% 96.3%
Table 4: Top-1 Accuracy in DVM-CAR Classification.
Next, we evaluate our KGV method in the DVM-CAR dataset. We see that even though the test distribution and training distribution have no significant data distribution shift, our KGV method still improves performance by incorporating multi-modal prior knowledge. The baselines we choose are contrastive learning approaches such as SimCLR (Chen et al. 2020) and SupCon (Khosla et al. 2020) as well as the SOTA model MMCL (Hager, Menten, and Rueckert 2023), which combines the tabular data and image data to enhance the model’s performance through contrastive learning. As shown in Table 4, our model outperforms all other baselines. In the mini-ImageNet domain, KGV achieves a 4.2% performance gain compared to the ResNet50 baseline. More ablation studies and insights can be found in the appendix – Additional Analysis section.
@@
Figure 3 shows the evaluation results of our KGV method in low data regimes. It illustrates that our KGV method significantly outperforms the baseline ResNet50, particularly when the amount of training data is limited. With the increase of training data, the performance of the two models gradually converges. However, our KGV method consistently outperforms the baseline among all training settings. The performance of KGV- – our model trained without synthetic images – is also shown in the figure. Without these visual priors, the performance of KGV drops.
@@
Figure 3: Performance under Low Data Regime. KGV- Indicates the KGV model trained without synthetic images.

### v2

# Robust Visual Representation Learning With Multi-modal Prior Knowledge For Image Classification Under Distribution Shift

@@ Abstract
Despite the remarkable success of deep neural networks (DNNs) in computer vision, they fail to remain high-performing when facing distribution shifts between training and testing data. In this paper, we propose Knowledge-Guided Visual representation learning (KGV) – a distribution-based learning approach leveraging multi-modal prior knowledge – to improve generalization under distribution shift. It integrates knowledge from two distinct modalities: 1) a knowledge graph (KG) with hierarchical and association relationships; and 2) generated synthetic images of visual elements semantically represented in the KG. The respective embeddings are generated from the given modalities in a common latent space, i.e., visual embeddings from original and synthetic images as well as knowledge graph embeddings (KGEs). These embeddings are aligned via a novel variant of translation-based KGE methods, where the node and relation embeddings of the KG are modeled as Gaussian distributions and translations, respectively. We claim that incorporating multi-model prior knowledge enables more regularized learning of image representations. Thus, the models are able to better generalize across different data distributions. We evaluate KGV on different image classification tasks with major or minor distribution shifts, namely road sign classification across datasets from Germany, China, and Russia, image classification with the mini-ImageNet dataset and its variants, as well as the DVM-CAR dataset. The results demonstrate that KGV consistently exhibits higher accuracy and data efficiency across all experiments.
Refer to captionRoad Sign Domain(a)Refer to captionGermany:(Train)China:(Test)Russia:(Test)Refer to captionCar DomainRefer to captionTrain:Test:Refer to captionImageNet DomainRefer to captionImageNet:(Train)ImageNet-V2:(Test)ImageNet-R:(Test)Refer to captionRefer to captionDE(b)TriangleExclamationWhiteRedhasShapehasSignLegendhasBackgroundC.hasBorderC.Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionCNDanger SignhasShapehasSignLegendBlackYellowRefer to captionRefer to captionhasBackgroundC.Refer to captionRefer to captionhasBorderC.Refer to captionRefer to captionSilverhasColorRefer to captionRefer to captionCoupeRefer to captionRefer to captionhasBodytypehasBodytypeAutohasGearType45hasViewpointNumber2hasDoorNumberhasDoorN.135hasViewpointManuellhasGearTypeThings(c)FoodBeverageBeerRefer to captionRefer to captionEatCornHotdogRefer to captionRefer to captionPlaceNatureCliff Coral Reef SpeciesEukar.AnimalBirdRefer to captionRefer to captionMammaFungus
Figure 1: (a) Examples of distribution shift in road sign, ImageNet, and Car Domains. (b) Decomposition of elements w.r.t. the road signs and cars. It represents the association relations between object categories and object category elements. (c) Examples of the abstract representation of hierarchical relations in the ImageNet domain.
@@ Introduction
Deep neural networks (DNNs) have demonstrated outstanding performance in computer vision (CV) tasks. They achieved remarkable success by leveraging vast amounts of data and computational power. However, these models often struggle with generalization, especially when the test data distribution deviates from the training data distribution (Liang, He, and Tan 2023). This happens as machine learning theory assumes that training and test data are identically distributed. Therefore, DNNs tend to overfit to the training domain. The situation worsens under low data regimes (Adadi 2021), as DNNs are more prone to overfitting when insufficient training data exists. Figure 1a) illustrates three distinct examples of distribution shifts between train and test data, in particular the Road Sign, ImageNet and Car domain. A DNN-based classifier applied to the testing domain will deteriorate in accuracy. This happens because the visual elements of the images vary between the train and test domains. For example, in Germany, warning signs have a white background, whereas in China, they use a yellow background. Similarly, ImageNet and Car differentiate between train and test data because of their style.
@@
First, we capture information about the structures of objects in a KG (Hogan et al. 2021) to inform and enhance the learning process. For instance, warning signs are composed of a white triangle with a red border plus one pictographic element (cf. Figure 1b). Road signs in different countries share the same elements, such as the shape or the pictographic element, or both. Also, the hierachical relations are demonstrated in Figure 1c. Combining this information, we construct a KG and develop a translation-based KGE method to map the nodes of the KG to embeddings in the latent space. An object category (e.g., ‘danger sign’) is naturally represented by its set of image vectors in the latent space. To align image and object categories, we represent the latter by a Gaussian distribution in the common latent space. Thus, the embeddings of object category elements and their relations to object categories indirectly influence the formation of categories in the latent space.
Second, for each object category element, we generate their respective synthetic images (if applicable) to augment the dataset. For instance, in the road sign domain, these elements are shape, color, and legend within the sign. These images are considered as visual prior knowledge of the object categories defined in the KG. Therefore, in this paper, the prior knowledge we used has two distinct modalities: 1) the factual knowledge across different data distributions represented in a KG, and 2) generated synthetic images of object category elements aligned with their semantic representation in the KG.
We evaluate the performance of our approach under distribution shifts and low data regimes in the context of image classification. To represent the prior knowledge, we construct three targeted KGs for the domains of Road Sign Recognition, Car Recognition, and ImageNet. Our experimental results show that KGV outperforms the image classification baselines in all experiments. Additionally, we found that our method is more data-efficient. Our main contributions are summarized as follows:
@@
  • We evaluate KGV on various image classification datasets, namely road sign datasets, mini-ImageNet and its variants, as well as the DVM-CAR datasets. The results illustrate a notable improvement under distribution shift: 4.4% and 4.1% on average in the ImageNet and Road Sign Recognition domains. Moreover, our method improves the SOTA results on the DVM-CAR dataset and shows significant improvements on data efficiency in low data regimes.
  • We analyse whether KGV can be adapted to the SOTA vision foundation models like CLIP and DINOv2. The results indicate that integrating KGV, CLIP and DINOv2 exhibit a better performance across various datasets. The code base is available at https://anonymous.4open.science/r/kgv-D68F/.
@@
Implicit prior knowledge in pre-trained models like CLIP (Radford et al. 2021) and DINOv2 (Oquab et al. 2024) can be leveraged by continuing to fine-tune them. Including explicit knowledge from KGs can also enhance model’s ability in visual or textual tasks (Monka, Halilaj, and Rettinger 2022b). Some approaches (Annervaz, Chowdhury, and Dukkipati 2018; Bosselut et al. 2019; Liu et al. 2020) leverage KGs to enhance language models with better reasoning ability. In the field of computer vision, KGs also play an important role. Li et al. (2023a) introduce a KG that includes inter-class relationships to the visual-language model, yielding a more effective classifier for downstream tasks. Some approaches  (Zareian, Karaman, and Chang 2020; Yu et al. 2021) build scene graphs to enhance their models’ performance. There also exist some approaches use extra information in the KG to perform zero-shot learning tasks (Wang, Ye, and Gupta 2018; Kampffmeyer et al. 2019; Roy et al. 2020) or few-shot learning tasks (Li et al. 2023c; Tseng et al. 2020).
@@
Our approach is designed end-to-end by adding the regularization loss and cross-entropy loss together to train the network. The regularization loss aligns the KG embedding and the visual embedding. The cross-entropy loss is used for the image classification. The architecture of KGV is depicted in Figure 2. Our model consists of three phases: the knowledge modeling phase, the training phase, and the inference phase. Here, we take the road sign domain as an example to describe our model.
@@
1. Knowledge Modelling PhaseRoad Sign DomainColorsRefer to captionRefer to captionRefer to captionShapesRefer to captionRefer to captionRefer to captionSign LegendsRefer to captionRefer to captionRefer to captionRoad SignsRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionVisual Knowledge GraphRoad SignWarningMandatory Danger
@@
Each image is provided as input into the image encoder to obtain the respective image embedding $\boldsymbol{z^{I}}\in\mathbb{R}^{d}$ in the latent space, where $d$ is the size of the dimension. The relations from KG connecting object categories, their elements and their respective images are encoded and stored in the relation lookup table, which is represented as a list of embeddings $[\boldsymbol{z_{0}^{r}},\boldsymbol{z_{1}^{r}},...,\boldsymbol{z_{N_{r}}^{r}}]$, where each $\boldsymbol{z_{i}^{r}}$ is the $i$-th relation vector embedding with the size of $d$. Note that the 0-th relation represents the inclusion relation demonstrated with purple dashed arrows as illustrated in Figure 2. The hierarchy of road signs is preserved in the KG. For instance, the images of the danger signs in Germany belong to the German danger sign category, and they also belong to the danger sign category, warning sign category, and road sign category. We use $\boldsymbol{z_{0}^{r}}$ to represent all these inclusion relations. In terms of the other $N_{r}$ relations, if one node $h$ has one relation with another node $t$, then all the sub-category nodes of $h$ have such relation with the node $t$. For example, all warning signs have a triangle shape, including their sub-categories like danger and danger animal. The node lookup table is represented as a list of embeddings $[\boldsymbol{z_{1}^{o}},\boldsymbol{z_{2}^{o}},...,\boldsymbol{z_{N_{o}}^{o}}]$, where each $\boldsymbol{z_{j}^{o}}$ denotes the $j$-th embedding. $N_{o}$ is the total number of nodes in the KG. The node embeddings are considered as Gaussian embeddings $\boldsymbol{z_{j}^{o}}=(\boldsymbol{\mu_{j}},\boldsymbol{\Sigma_{j}})$, where $\boldsymbol{\mu_{j}}$ and $\boldsymbol{\Sigma_{j}}$ are the mean and variance of the Gaussian distribution. Note that nodes, their relations, and respective images are embedded in the same latent space.
@@
where $\mathcal{N}(\boldsymbol{z^{I}}+\boldsymbol{z_{i}^{r}};\boldsymbol{\mu_{j}},\boldsymbol{\Sigma_{j}})$ denotes the possibility density of the vector $\boldsymbol{z^{I}}+\boldsymbol{z_{i}^{r}}$ under the Gaussian distribution with parameters of $\boldsymbol{\mu_{j}}$ and $\boldsymbol{\Sigma_{j}}$. Different score function designs are also compared in our experiments.
@@
where M is the mask defined in Equation (1). $\textbf{S}\in R^{(N_{r}+1)\times N_{o}}$ and $\textbf{S}_{ij}=-\log(s(\boldsymbol{z^{I}},\boldsymbol{z_{i}^{r}},\boldsymbol{z_{j}^{o}}))$, following the negative log likelihood form. $\epsilon$ is the threshold score difference between positive and negative triplets. Finally, the total loss is defined by:
@@
In this section, we first explain the datasets and the respective knowledge graphs used in our experiments. Then, we analyze the results of supervised image classification experiments under varying conditions, including data distribution shifts, low-data regimes, and few-shot learning scenarios. The experimental details are listed in the appendix section.
@@
The constructed knowledge graph contains 202 nodes and 5 relations. The 5 relations we include are ‘has the background color of’, ‘has the border color of’, ‘has the sign legend of’, ‘has the shape of’, ‘instance/subclass’.
@@
We leverage WordNet (Miller 1995; Fellbaum 2010), which is a lexical database containing nouns, verbs, adjectives, and adverbs of the English language structured into respective synsets and create the subset of Wordnet by extracting the respective synsets of each label from the mini-ImageNet dataset. These synsets are grouped based on the lexical domain they pertain to, e.g., animal, artifact, or food; thus, we get hierarchical relations between the classes in the mini-ImageNet from the WordNet. We also include two association relations: ‘has color’ and ‘has sketch’. We use the generated pure color images in the domain of road signs and the sketch images from ImageNet-Sketch. It contains 234 nodes and 3 relations: ‘subclass/instance of’, ‘has color’ and ‘has sketch’.
@@ Results
We conducted a number of supervised image classification experiments with major and minor distribution shifts and low data regimes. Further, we evaluated the model’s adaptability on few-shot learning experiments using data with altered distributions.
@@
Prior Know. Model GTSRB (43) CTSD (25) RTSD (36)
Imp. KG VP
ResNet50 98.9% (+0.4%) 71.5% (+3.3%) 72.8% (+7.3%)
CLIP+LP 77.2% (+22.1%) 73.5% (+1.3%) 60.0%(+20.1%)
DINOv2+LP 78.7% (+20.6%) 28.9% (+49.5%) 59.9% (+20.2%)
DGP 99.0 %(+0.3%) 72.0% (+2.8%) 73.5% (+6.6%)
GCNZ 99.1% (+0.2%) 71.1% (+3.7%) 74.9% (+5.2%)
KGV (ours) 99.3% 74.8% 80.1%
@@
Table 1: Road Sign Classification Accuracy. Evaluation results with the test data of GTSRB and the full data of CTSD and RTSD. Imp. refers to the implicit prior knowledge in pre-trained models, while KG indicates explicit prior knowledge from the knowledge graph. VP denotes the visual prior associated with object category elements.
@@
We train our KGV model and ResNet50 with the GTSRB dataset plus generated synthetic images from scratch. It is then evaluated with test data of GTSRB with 43 classes, the full data of CTSD with 25 shared classes, and the full data of RTSD with 36 shared classes. The shared classes here indicate those classes that both datasets contain. As shown in Table 1, all models’ accuracy drops in distribution shift scenarios, i.e., from Germany to China or Russia. By incorporating the multi-modal prior knowledge, our KGV reaches an obvious improvement compared to other baselines, such as DGP (Kampffmeyer et al. 2019) and GCNZ (Wang, Ye, and Gupta 2018). These two are ZSL baselines, which leverage additional information from KGs. Fine-tuning pretrained CLIP and DINOv2 with linear probing does not work well in road sign domain. ‘ResNet50+’ refers to the model trained with additional synthetic images and ‘KGV-’ denotes the model trained without synthetic images. As the experiments indicate, our KGV model surpasses the ResNet50 baseline by 3.3% in CTSD and 7.3% in RTSD. We also observe that the inclusion of synthetic images featuring elements of object categories, such as colors and shapes, enhances the generalization ability of the KGV models across various data distributions. On the other hand, it does not have a significant effect on the ResNet50’s performance. The precision and recall for a subset of classes in CTSD are presented in Table 2. As the table illustrates, the ResNet50 fails to classify specific classes, including danger signs, pedestrian crossing, etc. In contrast, the KGV models exhibit significant improvements in both precision and recall.
Datasets Samples ResNet50 KGNN ResNet50+ KGV CLIP+LP DINOv2+LP
Mini-ImageNet [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 62.6% +4.2% 66.2% +0.6% 59.6% +7.2% 66.8% 87.0% 94.9%
ImageNetV2 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 50.5% +2.2% 52.5% +0.2% 48.0% +4.7% 52.7% 87.5% 91.5%
ImageNet-R [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 25.5% +15.9% 30.0% +11.4% 36.9% +4.5% 41.4% 90.6% 86.1%
ImageNet-A [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 4.7% +2.1% 5.3% +1.5% 5.0% +1.8% 6.8% 53.8% 84.5%
@@
We also evaluate our method in the ImageNet domain. The baselines we choose are ResNet50, DINOv2, CLIP, and KGNN (Monka, Halilaj, and Rettinger 2022a). KGNN represents another approach to integrate prior knowledge into learning pipeline through pre-training. Our model outperforms the baselines KGNN and ResNet50 in all these datasets. As Table 3 shows, these baselines fail, particularly when the distribution shift increases (from Mini-ImageNet to ImageNetV2, ImageNet-R, and ImageNet-A). Our KGV suffers less when data distribution changes. We also observe that incorporating visual priors without the guidance of symbolic knowledge from the KG may negatively impact performance in scenarios where the difference between training and test distributions is minimal. Compared to ResNet50, ResNet50+’s performance decreases from 62.6% to 59.6% on mini-ImageNet and from 50.5% to 48.0% on ImageNetV2. We believe this is due to the visual priors altering the distribution of the training data, making it different from the test distribution. However, DINOv2 and CLIP fine-tuned with linear probing achieve better performance than our model, as they are pre-trained on massive data, including ImageNet and its variants. We conducted additional experiments by combining DINOv2 and CLIP with our approach, respectively, as described in Section Combining Implicit and Explicit Prior Knowledge.
Model SimCLR MMCL SupCon ResNet50 CLIP+LP DINOv2+LP KGV (Ours)
ACC 88.8% +7.5% 94.4% +1.9% 93.8% +2.5% 94.3% +2.0% 65.8% +30.5% 89.3% +7.0% 96.3%
Table 4: Accuracy in DVM-CAR Classification.
Next, we evaluate KGV in the DVM-CAR dataset. We observe that even though the test and training data have no significant distribution shift, it still improves performance. The baselines we choose are contrastive learning approaches such as SimCLR (Chen et al. 2020) and SupCon (Khosla et al. 2020) as well as the SOTA model MMCL (Hager, Menten, and Rueckert 2023), which combines the tabular data and image data to enhance the model’s performance through contrastive learning. Table 4 shows that our model outperforms all baselines. In the mini-ImageNet domain, KGV achieves a 4.2% performance gain compared to ResNet50.
@@
Figure 3 shows the evaluation results KGV in low data regimes. We see that our method significantly outperforms the ResNet50 baseline, particularly when the amount of training data is limited. With the increase of training data, the performance of the two models gradually converges. However, KGV consistently outperforms the baseline among all training settings. The performance of KGV- – the variant of KGV trained without synthetic images – is also shown in Figure 3. It indicates that without the visual priors, its performance drops.
@@
Figure 3: Performance under Low Data Regime. KGV- is the KGV variant trained without synthetic images.